OpenStudy (anonymous):

Algebra Two Help: A heated piece of metal cools according to the function c(x) = (.5)x − 10, where x is measured in hours. A device is added that aids in cooling according to the function h(x) = −x − 1. What will be the temperature of the metal after five hours? −6° Celsius 26° Celsius 32° Celsius 38° Celsius

OpenStudy (the_fizicx99):

is it \[C(x) = (0.5)^{x} - 10 \] Or, \[C(x) = (0.5)x -10\]

OpenStudy (anonymous):

[C(x) = (0.5)x -10\]

OpenStudy (the_fizicx99):

There's a difference, one has an exponent

OpenStudy (anonymous):

Ok, it is c(x)=(.5)^(x−10)

OpenStudy (the_fizicx99):

I'm mid way through it and something's wrong, is the 10 also an exponent?

OpenStudy (anonymous):

yes

OpenStudy (the_fizicx99):

Idk the wording is odd... The closest I got was -6 when I put -5-1 for c(x) I'm not sure if you specified correctly. >.<

OpenStudy (the_fizicx99):

@bibby Bestieeee idk they're asking for ;~;

OpenStudy (bibby):

my brain is not processing these words hold up You might have to add the functions

OpenStudy (anonymous):

I'm not sure but I think they are wanting me to somehow combine the two functions and then solve for c(5). I can solve for c(5), but I'm not sure how to combine them

OpenStudy (bibby):

(.5)x − 10 + −x − 1

OpenStudy (the_fizicx99):

I tried combining them and got a loong decimal, I think in c(x) only x is the exponent.

OpenStudy (the_fizicx99):

S(he) said there's an exponent 0.5^x-10 .____.

OpenStudy (bibby):

same thing. (.5)^(x − 10) + −x − 1

OpenStudy (anonymous):

fizi can you tell me how you made that .png file and I can show you the exact form the function is in

OpenStudy (the_fizicx99):

You can screen shot it, I use tool on the start tab

OpenStudy (bibby):

ms paint

OpenStudy (the_fizicx99):

It'd be -5 if the 10 is also an exponent, 0.5^5 = 0.03125 - 10 = -9.96875 .5^-5 = 32 Ohhh wait, its 26 degree's I think :O

OpenStudy (the_fizicx99):

-6 + 32 or 32-6 = 26 :3

OpenStudy (the_fizicx99):

-6 is from h(x) :>

OpenStudy (bibby):

I got the same thing doing it in my head. is that a choice?

OpenStudy (the_fizicx99):

it says 26 idk >.>

OpenStudy (bibby):

Choose B. 2 different people got it

OpenStudy (anonymous):

yeah 26 is a choice but how did you come to that?

OpenStudy (the_fizicx99):

Read above >.<

OpenStudy (bibby):

plug in 5 into:(.5)^(x − 10) + −x − 1 .5^-5 = 1/((.5)^5) + [-5-1] 1/((.5)^5) -6 etc.

OpenStudy (anonymous):

OH! Thank you guys so much!

OpenStudy (the_fizicx99):

:3

OpenStudy (bibby):

please be my internet girlfriend and give fiz the medal

OpenStudy (the_fizicx99):

What le fok >.>

OpenStudy (bibby):

owned

OpenStudy (anonymous):

I'm a guy...

OpenStudy (the_fizicx99):

:3 let's go gf

OpenStudy (bibby):

same but I identify as an anime girl

OpenStudy (the_fizicx99):

>.< bib les go

OpenStudy (the_fizicx99):

=.= do you actually have that on your computer?? >.> freak